Using the Node Voltage Method, what is the voltage v1 ? 2Ω 2 Ω 2Ω M 10V , 32Ω 2 Ω ΛΑ 6A
Using the Node Voltage Method, what is the voltage v1 ? 2Ω 2Ω 2 Ω 10V ν, 2 Ω Ο 2 Ω 6 Α.
Answer #1

2 Node voltage method => W+10 + V, Ye V, V, V 20. =v***-)-+) =>vi(+1) -V7( - Vi (2) – V2 (1) = 5 +0. Data + V2 - V1 + V-VI-68Ş-V, +V2 (1.5) = 6 . 5-V, + (2V, -5)(1.5) =6 , >-V, + 3V, -7.5=6 -5 2 V, = 13.5 => V, = 6.75V.